Logiweb dictionary of peano commutativity

Up Help

0 0 peano commutativity

1 0 M Lemma 1.8 2 0 M Proposition 3.2(a) 3 0 M Proposition 3.2(b) 4 0 M Proposition 3.2(c) 5 0 M Proposition 3.2(d) 6 0 M Proposition 3.2(f) (i) 7 0 M Proposition 3.2(f) (ii) 8 0 M Proposition 3.2(f) 9 0 M Proposition 3.2(g) (i) 10 0 M Proposition 3.2(g) (ii) 11 0 M Proposition 3.2(g) 12 0 M Proposition 3.2(h)(i) 13 0 M Proposition 3.2(h)_{(g)(ii)} 14 0 M Proposition 3.2(h)(ii) 15 0 M Proposition 3.2(h) 16 0 M Tautology A 17 0 M Tautology B

The pyk compiler, version 0.grue.20050603 by Klaus Grue GRD-2005-07-03.UTC:14:50:38.354331 = MJD-53554.TAI:14:51:10.354331 = LGT-4627119070354331e-6